OpenStudy (anonymous):

(x-2)^2/9+(y-4)^2/16=1 What are the major and minor axis

OpenStudy (mathmale):

The major axis is the longer one; the minor axis is the shorter one. You might find this problem easier to visualize and solve if you'd write it in standard format:\[\frac{ (x-2)^2 }{ 3^2 }+\frac{ (y-4)^2 }{ 4^2 }=1\]

OpenStudy (mathmale):

is the larger denominator quantity under the (x-2) or under the (y-4)? What is the significance of that?

OpenStudy (anonymous):

Sorry I didn't write in standard form, I can't figure out how to do that on here. I know that it is vertical but can't figure out if the major is y=4 or y-axis